# Machine Learning-based Near-field Emitter Location Sensing via Grouped Hybrid Analog and Digital XL-MIMO Receive Array

### v2

@@ I Introduction
NF localization is a fundamental technology in NF communication, and plays an important role in applications like channel estimation[4], integrated sensing and communication (ISAC) [5], physical layer security [6], etc.. The existing NF localization methods can be divided into two categories, the first is based on the conventional estimation methods. In [7], the conventional multiple signal classification (MUSIC) and maximum likelihood estimator (MLE) were first applied for the NF passive localization. The work in [8] proposed a two-stage MUSIC algorithm for localization of mixed NF and FF sources. A modified MUSIC algorithm for NF localization was proposed in [9], which has a reduced dimension compared to 2D-MUSIC algorithm and achieves much lower complexity without losing performance. The second category of algorithms are on the basis of high-order statistics. The work in [10] firstly proposed an NF localization algorithm by using fourth-order statistics. To further improve the performance, a second-order statistics-based weighted linear prediction algorithm was proposed in [11].
Recently, machine learning (ML) or deep learning (DL) - based methods have become popular in emitter localization because they can provide high estimation accuracy in both FF and NF scenarios. A deep neural network with a regression layer was adopted in [12] to address the problem of NF source localization. In [13], a complex ResNet was designed for NF DOA estimation. The work in [14] introduced a neural network for joint optimization of beamformer and localization function to achieve NF channel estimation. For the mixed NF and FF scenarios, the work in [15] developed convolution neural networks (CNN) for sources classification and localization by using the geometry of symmetric nested array. Classical clustering methods like $k$-means, density-based spatial clustering of applications with noise (DBSCAN) are also applied to FF DOA estimation[16]. Based on the existing works, ML-based methods can provide significant performance gains in the resolve of localization problems.
@@ II System Model
where $s(t)$ denotes the signal waveform, $v_{m}(t)\sim\mathcal{CN}(0,\sigma_{v}^{2})$ is the additive white Gaussian noise (AWGN), and $\varphi_{m}$ represents the phase difference caused by the signal propagation delay between the $m$th antenna and the reference point. Based on the uniform spherical model, $\varphi_{m}$ is given by
$\varphi_{m}=\frac{2\pi}{\lambda}\left(\sqrt{r^{2}+(m-1)^{2}d^{2}-2(m-1)dr\sin\theta}-r\right),$ (2)
@@ IV-B RSD-ASD-based Density Clustering
where $\bm{\omega}_{m,n}^{r}$ denotes the $n$-th elements of ~$\tilde{\mathcal{R}}_{imp,m}$ and $1\leq m\leq M_{imp}$. Due to the elements of true solution set have the tightest connections, so the value of $\epsilon$ is set as the minimum $d_{m,max}$, i.e.,
@@ V-A Network Construction
where $\hat{\bm{\Theta}}_{true}=[\hat{\theta}_{1,true},\cdots,\hat{\theta}_{L,true}]^{T}\in\mathbb{R}^{L\times 1}$ and $\hat{\bm{\Theta}}=[\hat{\theta}_{1,0},\cdots,\hat{\theta}_{1,M_{s}-1},\cdots,\hat{\theta}_{L,0},\cdots,\hat{\theta}_{L,M_{s}-1}]^{T}\in\mathbb{R}^{LM_{s}\times 1}$ are vectors that including the elements of sets $\hat{\Theta}_{true}$ and $\hat{{\Theta}}$, $\mathbf{W}^{(\rm{in})}$ and $\mathbf{b}^{(\rm{in})}$ are weight and bias vectors of input layer, $f^{(h)}(\cdot)$ represents the nonlinear transformation of the $h$th hidden layer. As the nonlinear modeling ability of neural networks comes from the activation functions, so here rectification linear unit (ReLU) is adopted as activation functions for the hidden layers. Finally, the expected output vector $\hat{\bm{\Theta}}_{true}$ contains $L$ specific angles, so this MLNN is designed for achieving a regression task, then the output layer can adopt linear activation function.
@@ V-B Training Strategy
The input of the MLNN is $\hat{\bm{\Theta}}$, and the expected output is $\hat{\bm{\Theta}}_{true}$, so the training data and label pair is given as $(\hat{\bm{\Theta}},\hat{\bm{\Theta}}_{true})$. And as the range of DOA is $[-\pi/2,\pi/2]$, different angles can generate different training data and label pairs, then the complete training set can be expressed by $\{(\hat{\bm{\Theta}}^{\theta},\hat{\bm{\Theta}}^{\theta}_{true})|\theta\in[-\pi/2,\pi/2]\}$. In the training stage, we choose mean square error (MSE) as the loss function, which is defined as
@@ VI-C CRLB Analysis
$\begin{cases}r_{l}\cos\theta_{l}=r\cos\theta_{l}\\ r_{l}\sin\theta_{l}=r\sin\theta-\Delta d_{l}\end{cases},$ (57)
@@
Lemma 1: The CRLBs for the DOA and range estimations of grouped hybrid architecture can be expressed by
@@ VII Simulation Results
In this section, we present some simulation results to evaluate the performances of the three proposed localization methods: MSDC, RSD-ASD-DBSCAN and RegNet. The CRLB of the proposed grouped hybrid structure is also considered as benchmark. Then the receive array employed in simulations is a ULA with $M=240$, $M_{s}=3$ and $L=5$, carrier frequency $f_{c}=30\rm{GHz}$, the emitter position is set as $(\theta,r)=(60^{\circ},20m)$. Under this condition, the MLNN part of the proposed RegNet is a fully-connected network has $LM_{s}=15$ input neurons, two hidden layers with 12 and 8 neurons respectively, and also has $L=5$ output neurons. The RegNet is trained for 100 epochs and optimized by Adam. The localization accuracy of the proposed methods are evaluated by the root mean-squared error (RMSE) of DOA and range estimations, and all results are averaged over 10000 Monte-Carlo simulations.
@@ Appendix A Derivation of CRLB
where $\bm{\beta}=[\theta,r]^{T}$, $1\leq i,j\leq 2$ and $F_{\beta_{i}\beta_{j}}^{(l)}=\sum_{l=1}^{L}{\rm{tr}}\left\{\mathbf{R}_{l}^{-1}\frac{\partial\mathbf{R}_{l}}{\partial\beta_{i}}\mathbf{R}_{l}^{-1}\frac{\partial\mathbf{R}_{l}}{\partial\beta_{j}}\right\}$. Then referring to (59), we can first attain two basic terms of $F_{\beta_{i}\beta_{j}}^{(l)}$

### v3

@@ I Introduction
NF localization is a fundamental technology in NF communication, and plays an important role in applications like channel estimation[4, 5, 6], integrated sensing and communication (ISAC) [7], physical layer security [8], etc.. The existing NF localization methods can be divided into two categories, the first is based on the conventional estimation methods. In [9], the conventional multiple signal classification (MUSIC) and maximum likelihood estimator (MLE) were first applied for the NF passive localization. The work in [10] proposed a two-stage MUSIC algorithm for localization of mixed NF and FF sources. A modified MUSIC algorithm for NF localization was proposed in [11], which has a reduced dimension compared to 2D-MUSIC algorithm and achieves much lower complexity without losing performance. The second category of algorithms are on the basis of high-order statistics. The work in [12] firstly proposed an NF localization algorithm by using fourth-order statistics. To further improve the performance, a second-order statistics-based weighted linear prediction algorithm was proposed in [13].
Recently, machine learning (ML) or deep learning (DL) - based methods have become popular in emitter localization because they can provide high estimation accuracy in both FF and NF scenarios. A deep neural network with a regression layer was adopted in [14] to address the problem of NF source localization. In [15], a complex ResNet was designed for NF DOA estimation. The work in [16] introduced a neural network for joint optimization of beamformer and localization function to achieve NF channel estimation. For the mixed NF and FF scenarios, the work in [17] developed convolution neural networks (CNN) for sources classification and localization by using the geometry of symmetric nested array. Classical clustering methods like $k$-means, density-based spatial clustering of applications with noise (DBSCAN), and global maximum cos_similarity clustering are also applied to FF DOA estimation[18, 19, 20]. Based on the existing works, ML-based methods can provide significant performance gains in the resolve of localization problems.
@@ II System Model
where $s(t)$ denotes the signal waveform, $v_{m}(t)\sim\mathcal{CN}(0,\sigma_{v}^{2})$ is the additive white Gaussian noise (AWGN), and $\varphi_{m}$ represents the phase difference caused by the signal propagation delay between the $m$th antenna and the reference point. Based on the uniform spherical model, $\phi_{m}$ is given by
$\phi_{m}=\frac{2\pi}{\lambda}\left(\sqrt{r^{2}+(m-1)^{2}d^{2}-2(m-1)dr\sin\theta}-r\right),$ (2)
@@ IV-B RSD-ASD-based Density Clustering
where $\bm{\omega}_{m,n}^{r}$ denotes the $n$-th sample of ~$\tilde{\mathcal{R}}_{imp,m}$ and $1\leq m\leq M_{imp}$. Due to the elements of true solution set have the tightest connections, so the value of $\epsilon$ is set as the minimum $d_{m,max}$, i.e.,
@@ V-A Network Construction
where $\hat{\bm{\Theta}}_{true}=[\hat{\theta}_{1,true},\cdots,\hat{\theta}_{L,true}]^{T}\in\mathbb{R}^{L\times 1}$ and $\hat{\bm{\Theta}}=[\hat{\theta}_{1,0},\cdots,\hat{\theta}_{1,M_{s}-1},\cdots,\hat{\theta}_{L,0},\cdots,\hat{\theta}_{L,M_{s}-1}]^{T}\in\mathbb{R}^{LM_{s}\times 1}$ are vectors that including the elements of sets $\hat{\Theta}_{true}$ and $\hat{{\Theta}}$, $\mathbf{W}^{(\rm{in})}$ and $\mathbf{b}^{(\rm{in})}$ are weight and bias vectors of input layer, $f^{(h)}(\cdot)$ represents the nonlinear transformation of the $h$th hidden layer. As the nonlinear modeling ability of neural networks comes from the activation functions, so here rectified linear unit (ReLU) is adopted as activation functions for the hidden layers. Finally, the expected output vector $\hat{\bm{\Theta}}_{true}$ contains $L$ specific angles, so this MLNN is designed for achieving a regression task, then the output layer can adopt linear activation function.
@@ V-B Training Strategy
The input of the MLNN is $\hat{\bm{\Theta}}$, and the expected output is $\hat{\bm{\Theta}}_{true}$, so the training data and label pair is given as $(\hat{\bm{\Theta}},\bm{\Theta}_{true})$. And as the range of DOA is $[-\pi/2,\pi/2]$, different angles can generate different training data and label pairs, then the complete training set can be expressed by $\{(\hat{\bm{\Theta}}^{\theta},\bm{\Theta}^{\theta}_{true})|\theta\in[-\pi/2,\pi/2]\}$. In the training stage, we choose mean square error (MSE) as the loss function, which is defined as
@@ VI-C CRLB Analysis
$\begin{cases}r_{l}\cos\theta_{l}=r\cos\theta\\ r_{l}\sin\theta_{l}=r\sin\theta-\Delta d_{l}\end{cases},$ (57)
@@
Lemma 1: The CRLBs for the DOA and range estimations of grouped HAD architecture can be expressed by
@@ VII Simulation Results
In this section, we present some simulation results to evaluate the performances of the three proposed localization methods: MSDC, RSD-ASD-DBSCAN and RegNet. The CRLB of the proposed grouped HAD structure is also considered as benchmark. Then the receive array employed in simulations is a ULA with $M=240$, $M_{s}=3$ and $L=5$, carrier frequency $f_{c}=30\rm{GHz}$, the emitter position is set as $(\theta,r)=(60^{\circ},20m)$. Under this condition, the MLNN part of the proposed RegNet is a fully-connected network has $LM_{s}=15$ input neurons, two hidden layers with 12 and 8 neurons respectively, and also has $L=5$ output neurons. The RegNet is trained for 100 epochs and optimized by Adam. The localization accuracy of the proposed methods are evaluated by the root mean-squared error (RMSE) of DOA and range estimations, and all results are averaged over 10000 Monte-Carlo simulations.
@@ Appendix A Derivation of CRLB
where $\bm{\beta}=[\theta,r]^{T}$, $1\leq i,j\leq 2$ and $F_{\beta_{i}\beta_{j}}^{(l)}={\rm{tr}}\left\{\mathbf{R}_{l}^{-1}\frac{\partial\mathbf{R}_{l}}{\partial\beta_{i}}\mathbf{R}_{l}^{-1}\frac{\partial\mathbf{R}_{l}}{\partial\beta_{j}}\right\}$. Then referring to (59), we can first attain two basic terms of $F_{\beta_{i}\beta_{j}}^{(l)}$